A fixed mass of gas is heated at constant pressure from 20 degree celcius to 90 degree celcius If its volume at 20 degree celcius is 200 cm^3 , calculate the volume at 90 degree celcius
1
Expert's answer
2020-07-06T15:39:47-0400

At constant pressure

"\frac{V_2}{V_1}=\frac{T_2}{T_1}"

Hence

"V_2=\frac{T_2}{T_1}V_1=\frac{90+273}{20+273}\times 200\:\rm cm^3=248\:\rm cm^3"